Woodworking Inc. is a private company. The custom build furniture. On June 1, 2021 Jill Wong signs a contract for Woodworking Inc. to build her a custom dining room set. She has to pay 15% on signed of the contract and will pay the balance on deliver on August 1, 2021. The selling price was $10,000 and the cost to build the furniture was $2,800.


1
Expert's answer
2022-03-14T13:23:05-0400

percentage gross profit mark up = gross profit/cost of sales

gross profit = sales - cost of sales

Gross profit = 10000 - 2800 = 7200

Gross profit mark-up = 7200/2800 = 257.14%